# Reply with Sticker: New Dataset and Model for Sticker Retrieval

### v3

@@ V-B Compared Methods
w/o knowledge 18.18 29.22 39.61 39.36
TABLE III: Experimental results (%) of different LLMs with or without intention.
@@
V-C Main Results
For text-to-image retrieval methods, they prioritize capturing semantic relationships between text and image content. However, since they are not explicitly designed for sticker retrieval scenarios, exhibiting inferior performance compared to our framework. In addition, an interesting observation is the overall superior performance of text-to-image retrieval methods compared to sticker retrieval baseline methods. This disparity can be attributed in part to the model design of SRS and MOD, where both models are devised to retrieve suitable stickers from a limited set of similar sticker candidates. Consequently, more attention is devoted to distinguishing local information among similar sticker expressions. In contrast, our dataset encompasses a more diverse range of stickers in real-world conversations, requiring the model to analyze and recognize more complex visual features. This demonstrates the advantage of the new dataset and the proposed framework in this work.
For Large language models, we design a prompt that integrates the current session to generate the sticker description for each session leveraging the in-context learning ability of LLMs [40, 41]. We then retrieve the appropriate sticker based on the generated sticker description and the combination of intent and sticker attributes using OpenAI’s LLM-based embedding model (text-embedding-ada-003) [42]. We can observe that LLM models perform better than sticker retrieval and text-to-image retrieval methods. This superior performance is attributed to the extensive parameterization and complex network architecture of LLMs, which greatly enhance their ability to understand and generate complex language and image descriptions. For P@1, P@3, and P@5, ChatGLM3, and Llama3 perform the best, while for mAP, Llama3 achieves the highest performance, reaching 39.76%. Compared to our IGSR method, all baseline models perform significantly worse. This further emphasizes the effectiveness of our approach in intent derivation, highlighting the critical role of intent as a bridging component in the process.
V-D Ablation Study
@@
We also conduct an ablation study on the use of knowledge and attributes. The evaluation results are shown in Table V-B. The performances of all ablation models are worse than those of the complete model under all metrics, which demonstrates the necessity of each component in our approach. Note that the removal of attributes ("w/o attribute") results in considerable performance degradation, indicating that utilizing attributes can make better learning of sticker representation in different sticker properties. Notably, by observing the performance of w/o attribute on P@5, we find that the impact of attributes is not as significant when a larger number of stickers are recalled. In addition, the removal of commonsense knowledge ( "w/o knowledge" ) sharply degrades performance, which verifies the importance of knowledge in understanding conversation context. It is worth noting that the absence of the intention tag ("w/o intention") leads to a more significant decline in performance, with the mAP and P@5 scores decreasing by 0.62% and 14.01%, respectively. This demonstrates the crucial role of the intention tag in improving the accuracy and relevance of sticker retrieval tasks, highlighting the importance of considering users’ intentions in retrieving appropriate stickers.
V-E Effect of Using Intention
@@
To analyze the impact of intention, We introduce the intention into various LLMs as mentioned in the baseline. The results are shown in Table V-B. "w/ intention" indicates that the large language model first predicts the intent of the response and then generates the reply based on the previous conversation context and the intent. "Sticker Response" and "Utterance Response" indicate whether the model generates a description of the sticker or generates a response for the user. Overall, the sticker description approach performs better than the text response approach, indicating that generating a sticker description can more directly highlight the key points of the response, thus improving the retrieval of the corresponding sticker. Furthermore, we can also find that predicting the intention first and then generating the response is less effective than directly generating the response for most LLMs. That is to say, an incorrect initial prediction of intention can lead to inappropriate responses, ultimately reducing performance. This suggests that merely introducing intention does not guarantee improved model performance. In contrast, our approach enables the model to simultaneously perform sticker retrieval and learn response intent, resulting in superior performance.
@@
V-F Effect of Number of Utterances
To examine and analyze the impact of the number of utterances over the performance of our proposed Int-RA framework, we conduct experiments by varying the number from 2 to 10 and demonstrate the results in Figure 6. We observe a similar trend across all evaluation metrics: mAP, P@1, P@3, and P@5. The results initially increase until the number of utterances reaches 6, after which they decrease as the number of utterances continues to increase. Two potential reasons may explain this phenomenon. Firstly, in limited contexts, the model can effectively capture features, resulting in improved performance as the amount of information increases. Secondly, the utility of utterance context may play a role. Utterances appearing too early before the sticker response may be irrelevant to the sticker and introduce unnecessary noise. In this dataset, it appears that 6 utterances are optimal.
TABLE IV: Experimental results (%) of different attributes. $\surd$ represents the used attribute. Ges., Pos., Face., and Ver. indicate gesture, posture, facial expression, and verbal, respectively.
@@ V-F Effect of Number of Utterances
V-G Effect of Using Different Attributes
In the process of attribute-aware sticker representation, four visual attributes are utilized in our proposed Int-RA to represent the key expression information of the sticker. This section examines the effectiveness of different attributes. The results of various scenarios of attribute combinations are shown in Table IV. It can be seen that the performance overall improves with the increase in the number of attributes used, using all the attributes achieves the best performance. Using only one attribute significantly degrades the performance, indicating that the visual information can not be learned sufficiently from a single perspective. That is, relying solely on this single attribute is insufficient for capturing the full expressive range necessary for accurate sticker representation. Further, incorporating multiple attributes provides richer information, thereby leading to improved performance. This concludes that a more holistic approach that combines multiple attributes to understand the visual information of stickers is essential for optimal performance.
Figure 7: Examples of conversation context and top-3 stickers retrieved by our method.
V-H Case Study
Several interactive cases retried by our approach are provided in Figure 7. These conversation samples suggest that our pipeline framework holds the capacity to provide sticker-incorporated expressive communication. From examples (a) and (b), we can observe that our approach tends to favor stickers with similar actions and facial expressions, with the characteristics of emoji stickers often being manifested through detailed features such as gestures and facial expressions. This demonstrates the effectiveness of using visual attributes to enhance the learning of stickers. However, in example (c), the final prediction is incorrect, likely due to the diverse styles of stickers, which remains a challenge in the current state of sticker recognition.
@@ V-H Case Study
V-I Visualization
To analyze how our Int-RA learns the important information about stickers, we visualize the relation score $\mathcal{M}$ (Equation 12) of three stickers in Figure 8. For example (a), where the character appears very angry. This indicates that the representation of this sticker heavily relies on this facial expression. Our Int-RA can effectively catch the important information in the sticker by the relation score placement on the character’s face. Moreover, the relation score also attends to the character’s gestures. For instance, in Case (b), where the character is depicted as holding chopsticks with one hand and supporting the face with the other, we observe attention focused on his hand, suggesting that our Int-RA learns key points of body language. Furthermore, considering that the relation score comprehensively considers four properties of stickers, as illustrated in Case (c), we observe that our Int-RA pays attention to both facial expressions and gestures simultaneously, thereby learning accurate visual information about the sticker.
@@ VI Conclusion and Future Work
We create a new dataset for sticker retrieval, called StickerInt. Unlike previous studies that view stickers merely as a supplement to the current utterance, our new dataset can cover two real-world scenarios of using stickers in online conversation: using stickers to reply to previous conversations or supplement our words. Based on the new dataset, we propose Int-RA, a framework for sticker retrieval in conversation. In which, the intention information is leveraged in the learning of conversation context. Further, we devise four novel visual attributes, i.e. gesture, posture, facial expression, and verbal, to improve the learning of stickers. Based on this, a relation-aware sticker selector is explored to retrieve the sticker for the conversation. Extensive experiments conducted on our StickerInt dataset demonstrate that our proposed approach achieves outstanding performance in sticker retrieval.
@@ VII Limitations
The limitations of this work are mainly twofold. Firstly, stickers have diverse styles ( e.g. cartoon, animal, etc.) in real-world conversations, which might affect the performance of the sticker retrieval task. Additionally, real-world conversations often involve multiple users engaging in multi-turn conversations. In such scenarios, our method may not fully capture the complexities of interactions among multiple users. Future research could focus on addressing these limitations by exploring more sophisticated models or incorporating additional contextual information to improve the performance of the sticker retrieval task.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62176076), the Natural Science Foundation of Guangdong (2023A1515012922), the Shenzhen Foundational Research Funding (JCYJ20220818102415032), and the Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies (2022B1212010005).

### v4

@@ V-D Results on MOD Dataset
-w/o knowledge 18.18 29.22 39.60 39.36
TABLE V: Experimental results (%) of different LLMs with or without intention.
@@
V-E Ablation Study
@@
We also conduct an ablation study on the use of knowledge and attributes. The evaluation results are shown in Table IV. The performances of all ablation models are worse than those of the complete model under all metrics, which demonstrates the necessity of each component in our approach. Note that the removal of attributes ("w/o attribute") results in considerable performance degradation, indicating that utilizing attributes can lead to better learning of sticker representation in different sticker properties. Notably, by observing the performance of w/o attribute on P@5, we find that the impact of attributes is not as significant when a larger number of stickers are recalled. In addition, the removal of the intention tag ( "w/o intention" ) sharply degrades performance, which verifies the importance of knowledge in understanding conversation context. This demonstrates the crucial role of the intention tag in improving the accuracy and relevance of sticker retrieval tasks, highlighting the importance of considering users’ intentions in retrieving appropriate stickers.
It is worth noting that the absence of the commonsense knowledge ("w/o knowledge") leads to a more significant decline in performance, with the mAP and P@3 scores decreasing by 14.01% and 8.44%, respectively. This quantitative result underscores that the commonsense knowledge derived from relations like xIntent, xNeed, xWant, xEffect, and xReact is not just a performance booster but a key component for the intention predictor to make more interpretable and contextually grounded predictions.
V-F Effect of Using Intention
@@
To analyze the impact of intention, we introduce the intention into various LLMs as mentioned in the baseline. The results are shown in Table V-D. "w/ intention" indicates that the large language model first predicts the intent of the response and then generates the reply based on the previous conversation context and the intent. "Sticker Response" and "Utterance Response" indicate whether the model generates a description of the sticker or generates a response for the user. Overall, the sticker description approach performs better than the text response approach, indicating that generating a sticker description can more directly highlight the key points of the response, thus improving the retrieval of the corresponding sticker. Furthermore, we can also find that predicting the intention first and then generating the response is less effective than directly generating the response for most LLMs. That is to say, an incorrect initial prediction of intention can lead to inappropriate responses, ultimately reducing performance. This suggests that merely introducing intention does not guarantee improved model performance. In contrast, our approach enables the model to simultaneously perform sticker retrieval and learn response intent, resulting in superior performance.
@@
V-G Effect of Number of Utterances
To examine and analyze the impact of the number of utterances on the performance of our proposed Int-RA framework, we conduct experiments by varying the number from 2 to 10 and demonstrate the results in Figure 6. We observe a similar trend across all evaluation metrics: mAP, P@1, P@3, and P@5. The results initially increase until the number of utterances reaches 6, after which they decrease as the number of utterances continues to increase. One potential reason may explain this phenomenon. While more context (up to six utterances) enriches the model’s understanding and improves performance, excessive context introduces noise from irrelevant earlier turns, impairing accuracy. Our findings show that a six-utterance context strikes the optimal balance between information richness and noise, leading to the best performance for sticker selection.
TABLE VI: Experimental results (%) of different attributes. $\surd$ represents the used attribute. Ges., Pos., Face., and Ver. indicate gesture, posture, facial expression, and verbal, respectively.
@@ V-G Effect of Number of Utterances
V-H Effect of Using Different Attributes
In the process of attribute-aware sticker representation, four visual attributes are utilized in our proposed Int-RA to represent the key expression information of the sticker. This section examines the effectiveness of different attributes. The results of various scenarios of attribute combinations are shown in Table VI. It can be seen that the performance overall improves with the increase in the number of attributes used, and using all the attributes achieves the best performance. Using only one attribute significantly degrades the performance, indicating that the visual information can not be learned sufficiently from a single perspective. That is, relying solely on this single attribute is insufficient for capturing the full expressive range necessary for accurate sticker representation. Further, incorporating multiple attributes provides richer information, thereby leading to improved performance. This concludes that a more holistic approach that combines multiple attributes to understand the visual information of stickers is essential for optimal performance.
Figure 7: Examples of conversation context, commonsense information, predicted intent, and top-3 stickers retrieved by our method.
V-I Discussion on Computational Efficiency and Scalability
To further explore the practicality of the proposed Int-RA framework, we examine its computational efficiency and scalability. The multi-stage architecture of Int-RA, comprising commonsense-based intention prediction, attribute-aware sticker description generation via MLLM, and relation-aware cross-modal matching, contributes to increased inference time. We analyze the overall computational cost of Int-RA in terms of three primary components: intention prediction time $t_{int}$, attribute-aware representation generation time $t_{attr}$, and relation-aware matching time $t_{match}$. The total inference time can be approximated as: $t=t_{int}+t_{attr}+t_{match}=O(n_{r}\cdot L)+O(k\cdot d)+O(k\cdot d^{2})$. where $n_{r}$ is the number of COMET relations, $L$ is the input sequence length for commonsense inference, $k$ is the number of candidate stickers, and $d$ is the embedding dimension for both visual and textual representations. In our experiment, Int-RA achieves an average response time of approximately 1.25 seconds on a single V100 GPU, which is acceptable for most real-time human-computer interaction applications.
Additionally, as demonstrated in our ablation studies (Section V-E), each component meaningfully contributes to final performance. Nevertheless, simplified variants of the model still retain competitive performance, offering flexibility in efficiency-performance trade-offs.
V-J Case Study
Several interactive cases retrieved by our approach are provided in Figure 7. These conversation samples suggest that our pipeline framework holds the capacity to provide sticker-incorporated expressive communication. From examples (a) and (b), we can observe that our approach tends to favor stickers with similar actions and facial expressions, with the characteristics of emoji stickers often being manifested through detailed features such as gestures and facial expressions. This demonstrates the effectiveness of using visual attributes to enhance the learning of stickers. However, in example (c), the final prediction is incorrect, likely due to the diverse styles of stickers, which remains a challenge in the current state of sticker recognition.
@@ V-J Case Study
To investigate the interpretability of the intention predictor, we analyze how commonsense knowledge contributes to intention inference. In example (a), the generated commonsense inferences include phrases such as “get rejected” and “feel awkward”, which strongly suggest a feeling of sadness. The predicted intention aligns with this emotional tone, enabling the model to select a sticker that appropriately conveys the underlying sentiment. This example illustrates how the intention predictor functions as a transparent and semantically grounded bridge between dialogue understanding and sticker selection. By leveraging external knowledge to infer user intent, the model enhances both retrieval accuracy and human interpretability.
@@
V-K Visualization
To analyze how our Int-RA learns the important information about stickers, we visualize the relation score $\mathcal{M}$ (Equation 12) of three stickers in Figure 8. For example (a), where the character appears neutral. This indicates that the representation of this sticker heavily relies on this facial expression. Our Int-RA can effectively catch the important information in the sticker by the relation score placement on the character’s face. Moreover, the relation score also attends to the character’s gestures. For instance, in example (b), where the character is depicted as holding chopsticks with one hand and supporting the face with the other, we observe attention focused on his hand, suggesting that our Int-RA learns key points of body language. Furthermore, considering that the relation score comprehensively considers four properties of stickers, as illustrated in example (c), we observe that our Int-RA pays attention to both facial expressions and gestures simultaneously, thereby learning accurate visual information about the sticker.
@@ VI Conclusion and Future Work
We create a new dataset for sticker retrieval, called StickerInt. Unlike previous studies that view stickers merely as a supplement to the current utterance, our new dataset can cover two real-world scenarios of using stickers in online conversation: using stickers to reply to previous conversations or supplement our words. Based on the new dataset, we propose Int-RA, a framework for sticker retrieval in conversation. In which, the intention information is leveraged in the learning of conversation context. Further, we devise four novel visual attributes, i.e., gesture, posture, facial expression, and verbal, to improve the learning of stickers. Based on this, a relation-aware sticker selector is explored to retrieve the sticker for the conversation. Extensive experiments conducted on our StickerInt dataset and other public datasets demonstrate that our proposed approach achieves outstanding performance in sticker retrieval.
@@ VII Limitations
The limitations of this work are mainly twofold. Firstly, stickers have diverse styles ( e.g., cartoon, animal, etc.) in real-world conversations, which might affect the performance of the sticker retrieval task. Additionally, real-world conversations often involve multiple users engaging in multi-turn conversations. In such scenarios, our method may not fully capture the complexities of interactions among multiple users. Future research could focus on addressing these limitations by exploring more sophisticated models or incorporating additional contextual information to improve the performance of the sticker retrieval task.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62176076), the Natural Science Foundation of Guangdong (2023A1515012922), the Shenzhen Foundational Research Funding (JCYJ20220818102415032), the Major Key Project of PCL2023A09, Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.